OpenStudy (anonymous):

I know the answer but will give ribbon to the 1st person to answer correctly. 8-c=c-8=[-c]/2=(2+22)=c*6 ...... what is c

OpenStudy (anonymous):

Medal not ribbon....come on guys its easy why not reply

OpenStudy (anonymous):

you give up @UditKulka

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

what's c

OpenStudy (anonymous):

1 min i think i got it

OpenStudy (anonymous):

c=4?

OpenStudy (anonymous):

@awesomelyqwerty

OpenStudy (anonymous):

8-4=4-8=[-4]/2=(2+22)=4*6 which equals 24.... congratulations...lol...:)

OpenStudy (anonymous):

Thank uu

OpenStudy (anonymous):

your welcome .....you should give me a ribbon....just sayin

OpenStudy (anonymous):

But do u have any sophisticated method for doing this sum bcoz my method was weird

OpenStudy (anonymous):

@UditKulka yeah c-8=-c so c would have had to be the same number which was 4 then c/2=2 the only mathematical explanation had to be 4......

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Hey bro which grade r u in and from where u are

OpenStudy (anonymous):

9th and i was born in germany but raised in the U.S..... what about you

OpenStudy (anonymous):

i m in 5th i m an indian

OpenStudy (anonymous):

shouldnt you be in school...... or are you homeschooled

OpenStudy (anonymous):

yes i am in school

OpenStudy (anonymous):

hey but there is time difference in us and india so i m at home right now!

OpenStudy (anonymous):

oh....cool